OpenStudy (anonymous):

for which equation is x=-1/2 a solution? a. 7-4x=-9 b. 9x+2=7 c. 6+8x=-10 d. 10x-3=-8

OpenStudy (anonymous):

d. 10x -3 = -8 10x = -5 x = -5/10 x = -1/2